OpenStudy (anonymous):

Solve 5x + 5 = 9x Will give medal

OpenStudy (anonymous):

Wait the question is this

OpenStudy (ja1):

Yes we need to isolate the variable

OpenStudy (anonymous):



OpenStudy (abb0t):

take the natural log of both sides.

OpenStudy (abb0t):

use your log rules to isolate the "x". for instance: \(ln(a)^b = b~ln(a)\) and \(ln\frac{a}{b} = ln(a) - ln(b)\)

OpenStudy (anonymous):

\[5^{x+5}=9^{x}\]

OpenStudy (anonymous):

Ok, keep going